### v2

# NOLO: Navigate Only Look Once

Abstract
The in-context learning ability of Transformer models has brought new possibilities to visual navigation. In this paper, we focus on the video navigation setting, where an in-context navigation policy needs to be learned purely from videos in an offline manner, without access to the actual environment. For this setting, we propose Navigate Only Look Once (NOLO), a method for learning a navigation policy that possesses the in-context ability and adapts to new scenes by taking corresponding context videos as input without finetuning or re-training. To enable learning from videos, we first propose a pseudo action labeling procedure using optical flow to recover the action label from egocentric videos. Then, offline reinforcement learning is applied to learn the navigation policy. Through extensive experiments on different scenes both in simulation and the real world, we show that our algorithm outperforms baselines by a large margin, which demonstrates the in-context learning ability of the learned policy. For videos and more information, visit our project page.
1 Introduction
[…]
Visual navigation has long become a research focus due to its wide application in different fields, including mobile robots [18], autonomous vehicles [38], and virtual assistants [22]. Existing approaches to visual navigation can be broadly divided into two main categories: modular and end-to-end methods. Modular methods decompose the navigation problem into specific tasks, such as simultaneous localization and mapping, object segmentation, keypoint matching, and depth estimation [15, 16, 14, 91, 45]. End-to-end methods [37, 77, 10], on the other hand, bypass such modularization, directly learning the mappings from raw sensory observations (e.g., GPS and Compass [31, 65], RGB-D cameras [34], and IMU [58]) to action outputs. While these approaches have yielded promising results, they still face significant challenges, particularly in terms of generalizability. Modular methods can be brittle when transferred to unfamiliar environments. End-to-end approaches mainly focus on online learning [16, 88, 17], which requires continuous, interactive engagement with simulated or real environments. Thus, they may require additional interactions or exploration to adapt to novel scenes, complicating their deployment in real-world navigation settings.
To address these limitations, it is helpful to consider how humans intuitively navigate. We can often find our way through unfamiliar spaces after watching a simple traversal video. Our navigation abilities do not depend on precise positional data or distance measurement. Instead, we rely primarily on visual recognition, understanding, and memory. Building on this insight, we introduce a new visual navigation setting called Video Navigation, which mirrors how humans rely on visual observations to navigate. In video navigation, a mobile agent is provided with a context video of an unfamiliar environment and learns to reach any target seen in the video. For instance, a sweeping robot equipped with video navigation capabilities could immediately navigate and clean any area within a house after viewing a short video of its layout. We believe video navigation has immense potential for endowing mobile agents with human-like navigation capabilities. Although encouraging, achieving robust video navigation poses considerable challenges:
  • Limited Observations. Learning to navigate from egocentric traversal videos is challenging due to a lack of information, including actual actions, spatial cues, scene structure, and visual context.
  • Absence of Explicit Intent. Egocentric videos lack clear navigational intentions, making conventional methods like inverse reinforcement learning unsuitable, as these approaches typically require expert trajectory data.
  • Minimal Sensory Input. Camera intrinsics, robot poses, maps, odometers, or depth inputs are not necessary for video navigation, which distinguishes our work from others. A majority of existing methods prefer to build explicit spatial maps for planning with heterogeneous sensors. In this work, we aim to enable human-like navigation using only egocentric RGB observations.
Fundamentally, the objective of video navigation is to train a meta-policy that enables a robot to adapt to a novel scene–essentially, to navigate anywhere using cues from a context video. Recent research demonstrates that in-context learning (ICL [23]) excels at solving tasks using a few contextual demonstrations without requiring additional training [7] and has proven valuable across fields from LLM-based agents [85] to meta-learning [48, 49, 55]. Inspired by the success of ICL, we propose Navigate Only Look Once (NOLO), which learns a transferable in-context navigation policy conditioned on a context video, enabling adaptation to new scenes without fine-tuning or further re-training.
To achieve this, we collect a dataset of egocentric traversal videos, each covering a unique scene. We derive pseudo-actions from these context videos using optical flow, generating full frame-action trajectories, and then employ an offline reinforcement learning approach to train the navigation policy. This resource-efficient offline training pipeline eliminates the need for time-consuming online interactions, enabling large-scale training on simulated data and facilitating the transfer of navigation skills to unfamiliar environments. Furthermore, NOLO’s in-context adaptability makes it particularly suitable for real-world deployment, where frequent policy adjustments are often impractical.
Our key contributions can be summarized as follows:
  • We present a novel and practical setting Video Navigation, which necessitates learning an in-context navigation policy in an offline manner purely from videos such that the learned policy can adapt to different scenes by taking corresponding videos as context.
  • We propose NOLO to solve the video navigation problem. NOLO seamlessly incorporates optical flow into offline reinforcement learning via pseudo-action labeling. To encourage the temporal coherence of representation, we further propose a temporal coherence loss for temporally aligned context visual representation.
  • Empirical evaluations on RoboTHOR and Habitat benchmarks demonstrate the advantages of NOLO over baselines in video navigation. Notably, navigation is achieved by only viewing a single 30-second video clip per scene. We further deploy NOLO on the Unitree Go2 robot and experiments verify that NOLO is also effective and friendly to real-world deployment.
2 Related Work
Visual Navigation. The visual navigation problem can be categorized into indoor navigation [11, 33, 93, 30, 54, 61, 46, 77, 60, 42, 51] and outdoor navigation [52, 76, 86]. In this paper, we only focus on the indoor navigation problem. Current research mainly focuses on finding the location of specific positions [65, 41], rooms [62], or object instances [44, 45, 41, 50]. The goals of navigation can be described in coordinates [65], natural language [68, 91, 90, 10, 94], goal image of target position [62, 91, 90], or goal image of object instances [44, 45, 41, 50, 43, 60, 46]. Our video navigation of finding objects occurred in the context video falls into the last group. Visual navigation methods can also be categorized by whether exploration is required. Most existing works aim for generalization across new goals or environments, heavily relying on learning algorithms that require extensive online interactions, often through reinforcement learning [99, 80, 2], paired with real-time visual feedback [61]. This approach enables the agent to adapt dynamically but requires significant interaction within the environment. Some studies attempt to improve data efficiency but still rely on random policy exploration or assume access to rich datasets, including robot poses [14] or depth images [34, 30, 42, 45]. In contrast, NOLO learns entirely offline. It leverages video data for navigation, freeing the agent from real-world exploration and allowing effective policy learning without performing actual actions, marking a novel approach in video-based visual navigation.
Learning from Videos. There are four main lines of research study learning from videos. (1) Learn distinctive representations for intrinsic rewards. These approaches mainly include predicting the future [98, 25, 96], learning temporal abstractions [96, 75, 4, 8], contrastive learning [47, 79], adversarial learning [66, 82, 40, 57] or other transition-learning based methods [6, 53, 64]. (2) Recover value function along with representations [26, 59, 32, 5]. (3) Pretrain action-free model in the first stage and then boost the model with additional action information [92, 74, 87]. (4) Discover latent actions [24, 56, 72, 9] and then align to the real actions online in the environment [24, 56, 72], by human annotation [9] or from offline datasets [72]. Unlike previous work, we focus on the visual navigation problem and by utilizing its specificity, propose an offline method that directly learns the actual navigation policy, not the latent one, from videos. Few attempts are devoted to decoding real actions directly from videos. A recent work AVDC [43] attempts to principally use similar rule-based methods to directly infer actions for robot manipulation. However, it is an online algorithm with additional depth image inputs. We also compare NOLO with AVDC [43] in Sec. 4.5.
In-Context Learning. In-context learning (ICL), first demonstrated in large language models [7, 83, 19], enables models to perform new tasks with a few examples as context. This capability has extended to vision-language models [3, 84, 1], AI agents [78, 36], and models across heterogeneous modalities [39, 27]. Our work uniquely applies ICL to video navigation. To our knowledge, NOLO is the first attempt to learn an adaptable policy guided by context videos for diverse scenes.
3 Learning to Navigate from Videos
3.1 The Problem Formulation of Video Navigation
In video navigation, we offline recover a navigation policy $\pi_{\theta}\left(a|g,o,\mathcal{V}\right)$ to find objects occurred in the context video, which takes a goal image $g$, current observation $o$, and context video $\mathcal{V}$ as input and outputs discrete action $a$, considering high-level discrete actions are commonly used in competitions like the Habitat Navigation Challenge [69] and modern navigation tasks [52]. To train such a policy, we assume a video dataset $\mathcal{D}$ contains $N$ videos collected (e.g., by some simple rule-based policy or humans) in $N$ different scenes, $\mathcal{D}=\left\{\mathcal{V}^{1},\mathcal{V}^{2},...,\mathcal{V}^{N}\right\}$. Each video $\mathcal{V}^{i}$ contains $T$ frames $f^{i}_{t}$, $\mathcal{V}^{i}=\left\{f^{i}_{1},f^{i}_{2},...,f^{i}_{T}\right\}$. During inference, we deploy the learned policy in a new scene given a context video $\mathcal{V^{\textrm{new}}}\notin\mathcal{D}$ and a goal frame $g\in\mathcal{G}^{\textrm{new}}$ describing an object occurred in $\mathcal{V^{\textrm{new}}}$. The policy is expected to adapt to this new scene via in-context learning.
3.2 Pseudo Action Labeling
In video navigation, capturing environmental dynamics is crucial for effective adaptation to new scenes. Building on prior work [29, 24, 56, 72, 9] that leverages inverse dynamics models (IDMs) to extract latent actions from adjacent frames, we make a further step by converting latent actions into actual discrete actions within video navigation. Specifically, we propose pseudo action labeling, as illustrated in  Fig. 2. we employ GMFlow[89] as an implicit IDM $F_{\xi}$ to estimate discrete pixel displacements between consecutive frames and predict real navigation actions from flow maps with rule-based post-processing. For each context video $\mathcal{V}$, we label a sequence of pseudo-actions $\boldsymbol{\hat{a}}$ to get a trajectory $\mathcal{T}=\{f_{t},\hat{a}_{t}\}^{T}_{t=1}$. See Appendix 6.1 for more details.
Figure 2: Two adjacent frames are taken by a pretrained optical flow model to get a flow map. Some representative dominant vectors are filtered for action selection.
3.3 In-Context Policy Modeling
[…]
$h_{t+1},\pi_{\theta}\left({a}_{t}|g,o_{t},\mathcal{T}\right)=\text{VN}\circlearrowleft\text{Bert}(g,o_{t},\mathcal{T}),$ (1)
where $h$ is specially designed as a recurrent hidden state to capture informative attention history.
[…]
$\begin{split}\mathcal{L}_{a}=-\mathbb{E}_{g,o_{t},\hat{a}_{t}\sim\mathcal{T}^{i}}\log\pi_{\theta}\left(\hat{a}_{t}|g,o_{t},\mathcal{T}^{i}\right)\\ \mathcal{L}_{d}=-\mathbb{E}_{g,o_{t}\sim\mathcal{T}^{i}}\log\delta_{\upsilon}\left(d_{t}|g,o_{t},\mathcal{T}^{i}\right).\end{split}$ (2)
3.4 Batch-Constrained Q-Learning
Since the policy used to record the video may be highly suboptimal for video navigation, we favor offline reinforcement learning over imitation learning for policy learning. Specifically, we adopt BCQ [28], an offline reinforcement learning method that emphasizes constraining the action selection to those within the distribution of the observed frame-action pairs. Q-function $Q_{\omega}$ is learned from Bellman update in Eq. 3, where $r(g,o_{t},\hat{a}_{t})=\mathbbm{I}(o_{t}=g)$ is a binary reward indicating the success of a task,
$\begin{split}\mathcal{L}_{q}=\mathop{\mathbb{E}}_{o_{t},o_{t+1},g,\hat{a}_{t}\sim\mathcal{T}^{i}}\left[Q_{\omega}\left(g,o_{t},\hat{a}_{t},\mathcal{T}^{i}\right)-y_{t}\right]^{2},\\ y_{t}=r(g,o_{t},\hat{a}_{t})+\gamma\max_{\tilde{a}_{t+1}\in A^{\beta}}Q_{\omega}\left(g,o_{t+1},\tilde{a}_{t+1},\mathcal{T}^{i}\right).\end{split}$ (3)
Given the action space is discrete, we define a set $A^{\beta}$ in Eq. 4 to select $\tilde{a}_{t+1}$, where $\beta$ represents the threshold for action selection. Action choices are constrained based on the action distribution within the context trajectory dataset. The learned Q-function, $Q_{\omega}$, thus can wisely guides the agent toward the target goal $g$.
$A^{\beta}:=\Big{\{}\tilde{a}_{t+1}\Big{|}\frac{\pi_{\theta}\left(\cdot|g,o_{t+1},\mathcal{T}^{i}\right)}{\max\pi_{\theta}\left(\cdot|g,o_{t+1},\mathcal{T}^{i}\right)}>\beta\Big{\}}.$ (4)
3.5 Temporal Coherence
Learning temporally aligned representation of videos is crucial for understanding the causality of the context video. Inspired by RLHF [63], we learn a ”fuzzy” measure of progress from their natural temporal ordering, where the higher preference score is assigned to the later frames in the video. To achieve this, we learn a temporal indicator $\hat{u}_{\phi}(e)$, which consists of 1D self-attention and spectral normalization modules to map a visual context embedding $e^{s}_{t}$ to a utility score. $\hat{u}_{\phi}$ is optimized according to Eq. 5:
$\min_{\zeta,\phi}\mathcal{L}_{t}=\underset{f_{t_{-}},f_{t_{+}}\sim\mathcal{V}}{\mathbb{E}}-\log\sigma\left[\hat{u}_{\phi}\left(E_{\zeta}(f_{t_{+}})\right)-\hat{u}_{\phi}\left(E_{\zeta}(f_{t_{-}})\right)\right],$ (5)
where $t_{+}>t_{-}$ and $f_{t_{+}},f_{t_{-}}$ stands for the later and earlier frame in the video. This self-supervised paradigm naturally drives the temporal indicator $\hat{u}_{\phi}(s)$ to maximize likelihood to ensure the temporal order preference $f_{t_{+}}\succ f_{t_{-}}$, leading to temporally aligned context visual representation, which is beneficial to better understand the temporal relation in a novel scene, as elaborated in later Sec. 4.5.
4 Experiments
4.1 Experimental Setups
Configurations. We apply our algorithm to realistic embodied navigation testbeds RoboTHOR [20] and Habitat [71]. For each navigation task in RoboTHOR and Habitat, we initialize a mobile robot with a random position, orientation, and a given goal image extracted from the context video indicating a specific goal object. At each timestep, the agent, mounted with only one RGB camera of $224\times 224$ resolution, receives an egocentric image and can execute a discrete action from a fixed action space $[{\mathrm{MoveForward,TurnLeft,TurnRight,STOP}}]$. The minimum units of rotation and forward movement are 30° and 0.25 meters respectively. When the goal object is observed and the distance between the goal object and the agent is within 1m, the agent succeeds, otherwise it fails when the maximum step is reached, which is 500 in all experiments.
Task Division. In RoboTHOR, there are 15 different rooms and 5 possible layouts for each room, in total 75 different scenes. Each room has a different topology from other rooms, and the location of objects in each layout is different from other layouts. In order to test the in-context learning ability at different levels, we divide the training and test scenes as follows. For the training set, we select 12 rooms, and 4 layouts for each room. For the unseen layout testing set, we select the remaining 1 layout of the 12 training set rooms, in total 12 scenes. For the unseen room testing set, we select the remaining 3 rooms and all their layouts, in total 15 scenes. In Habitat, we divide all 90 building-scale scenes from Matterport3D (MP3D)  [12] into $70$ training scenes and $20$ testing scenes to test the generalization abilities. See Appendix 7.1 for details.
Dataset Construction. We use a fixed non-explorative rule-based policy to collect the video dataset. The policy keeps moving forward until colliding with an obstacle and then randomly turns to get rid of it. Simultaneously, the agent keeps recording an egocentric RGB sequence till reaching the maximum 900 steps, i.e. a 30-second video clip in each scene. After that, we employ a pretrained object detector Detic [97] to detect all the objects occurred in the video. To get a set for goal frames, we additionally roll out in the simulation to collect goal images of each target object from different views.
Task Evaluation. To evaluate the comprehensive ability of the in-context navigation policy, we design three levels of generalization test:
  • Evaluate in the unseen layout testing set to test the generalization ability over layouts.
  • Evaluate in the unseen room testing set to test the generalization ability over rooms.
  • Evaluate in scenes from a different simulator to test the generalization ability over domains.
For evaluation, the agent is arranged to navigate to each target object detected in a context video. Each navigation task repeats 10 times.
We conduct 3 runs with different random seeds and quantitatively assess the performance using metrics including success rate (SR), success weighted by normalized inverse path length (SPL), trajectory length (TL), and navigation error (NE) following a recent work [93].
Baselines. NOLO stands as the first method to incorporate context videos directly into policy for in-context navigation tasks. To address in-context generalization capabilities, we compare NOLO with a random exploration policy, as well as two large multimodal models (LMMs), GPT-4 and Video-LLaVA which also exhibit remarkable zero-shot multimodal understanding and generation abilities, in a training-free manner. To adapt the LMMs for video-based navigation, we use a queue of size $K$ to store context video frames, current observations, and goal images, which are concatenated as LMM input. A unified prompt guides the LMMs to understand the context and navigate effectively in both RoboTHOR and Habitat, extracting actions from their language responses. See Appendix 7.2 for details.
In addition, we further compare NOLO with traditional visual navigation methods without context video input, specifically, an online method VGM [46], which builds a memory graph based on unsupervised image representations from navigation history to guide actions, and an offline method ZSON [60], which encodes goal images into a multimodal, semantic embedding space to enable scalable training of SemanticNav agents in unannotated 3D environments. Besides, we compare NOLO with a recent work AVDC [43], which infers actions from video predictions while incorporating extra depth information. Our experimental results show that AVDC underperforms NOLO even with additional depth images.
4.2 RoboTHOR
We average the metrics of video navigation tasks in all testing scenes and showcase the mean results in Tab. 1. NOLO demonstrates overwhelming advantages over baselines, including random, LMM baselines, and visual navigation baselines in both testing sets, which proves its strong generalization abilities to new scenes in different generalization levels. The set of unseen rooms is naturally a harder testing set than the unseen layouts in terms of generalization ability. Therefore, NOLO displays slightly better navigation performance in the unseen layout testing set (71.92% SR, 29.26% SPL) compared to the unseen room testing set (70.48% SR, 27.74% SPL). GPT-4o outperforms the random exploration policy across most scenarios, while Video-LLaVA generally performs below the random exploration policy due to its smaller model size relative to GPT-4o. NOLO is also much more parameter-efficient than GPT-4o. However, it achieves better navigation results, underscoring its effectiveness in training in-context navigation policies. Overall, visual navigation baselines tend to surpass both random and LMM baselines because of additional training. However, AVDC exhibits instability in unseen environments due to its reliance on a diffusion model trained on a limited set of egocentric video frames, which leads to inconsistency in video prediction. While VGM’s memory graph and ZSON’s multimodal semantic embeddings can provide some guidance, they still fall short of NOLO in terms of success rate and path efficiency. We analyzed that VGM’s implicit graph struggles to effectively represent topological structures with limited data coverage, and for ZSON, there is a non-negligible distribution discrepancy between CLIP [67] pre-training data and our egocentric videos. See Figs. 9 and 10 in Appendix 9 for detailed results.
Table 1: Performance comparison of success rates (SR), success path lengths (SPL), trajectory lengths (TL), and navigation errors (NE) across different methods and environments. The blue part shows some training-free methods and the green part shows visual navigation approaches which require training.
Method Robothor Habitat
Unseen Layout Unseen Room Unseen Room
[…]
Random 26.00 13.27 394.77 2.93 23.10 12.02 405.87 2.89 24.53 10.03 402.02 4.97
GPT-4o 31.97 17.18 363.23 3.03 31.69 16.49 364.63 2.89 35.16 20.39 345.84 4.70
Video-LLaVA 20.07 8.31 381.92 3.08 17.51 6.58 385.72 3.32 14.55 9.02 384.53 4.71
AVDC 32.53 9.43 430.77 2.15 31.14 11.57 424.56 2.13 27.29 12.24 418.12 4.30
VGM 47.04 23.90 376.92 2.44 40.51 19.99 400.90 2.38 34.34 17.30 393.92 4.19
ZSON 37.83 21.31 349.31 2.23 35.78 18.81 372.68 2.37 32.83 19.86 395.94 4.32
NOLO (ours) 71.92 29.26 238.92 1.79 70.48 27.74 248.44 1.87 43.65 20.77 347.57 3.67
(a) H2R-SR
(b) H2R-SPL
(c) R2H-SR
(d) R2H-SPL
Figure 4: Average SR and SPL of cross domain evaluation. H2R means training in Habitat and testing in RoboTHOR, and vice versa.
Table 2: Ablation studies of NOLO. For each method, we average the metrics over all testing scenes and average all the mean results over three random seeds to report the mean and standard deviation.
Method Robothor Habitat
[…]
NOLO 71.20 $\pm$ 0.85 28.50 $\pm$ 0.72 243.68 $\pm$ 3.80 1.83 $\pm$ 0.04 43.65 $\pm$ 1.09 20.77 $\pm$ 0.46 347.57 $\pm$ 3.32 3.67 $\pm$ 0.09
NOLO-C 67.78 $\pm$ 0.60 27.70 $\pm$ 1.35 259.65 $\pm$ 0.53 1.98 $\pm$ 0.02 33.58 $\pm$ 0.92 17.41 $\pm$ 0.43 381.26 $\pm$ 3.23 4.58 $\pm$ 0.26
NOLO-T 69.03 $\pm$ 0.87 28.15 $\pm$ 0.09 252.84 $\pm$ 6.33 1.93 $\pm$ 0.04 37.48 $\pm$ 1.01 18.51 $\pm$ 0.71 369.85 $\pm$ 2.72 4.72 $\pm$ 0.07
NOLO(M) 66.15 $\pm$ 0.37 26.66 $\pm$ 1.61 267.95 $\pm$ 1.96 2.02 $\pm$ 0.00 36.73 $\pm$ 1.89 18.17 $\pm$ 0.31 373.92 $\pm$ 3.46 4.69 $\pm$ 0.25
Figure 5: Visualizations of six video navigation tasks during in-context policy deployment. The leftmost bird-eye-view topological maps depicts the robot positions corresponding to the key observations to the right, and the second column displays the goal images.
4.3 Habitat
We evaluate NOLO similarly in Habitat to demonstrate the effectiveness of NOLO in larger-scale scenes. As illustrated in Tab. 1 , NOLO still outperforms baselines in terms of SR and achieves slightly better performance in terms of SPL. Compared with the metrics of NOLO in RoboTHOR, the average SPL and SR are lower, and the average NE and TL are higher. This is mostly because scenes in Habitat are significantly larger than those in RoboTHOR, resulting in more challenging navigation tasks across multiple rooms and even floors and requiring longer traversal distances. However, we observe that NOLO is still capable of acquiring many necessary sub-skills like entering or exiting a room, avoiding obstacles, and exploring. See Appendix 9 for detailed results.
4.4 Cross-Domain Evaluation
To pursue the ultimate goal of video navigation, we evaluate NOLO across RoboTHOR and Habitat, even if the visual inputs may suffer from a severe domain gap. Specifically, NOLO trained from 48 RoboTHOR scenes is evaluated in 20 Habitat scenes (R2H), and NOLO trained from 70 Habitat scenes is evaluated in 27 RoboTHOR scenes (H2R). For comparison, we also plot the average in-domain results (R2R and H2H). We average all the results on testing scenes in Fig. 4, from which we surprisingly discover that R2H and H2R show competitive performance compared to R2R and H2H. This is because the pretrained image encoder in NOLO provides a general representation of indoor scenes, and the temporal coherence loss in Eq. 5 also leads to a general representation reflecting the natural temporal ordering, which together lead to the generalization ability of NOLO across domains.
4.5 Ablation Studies
In this section, we use a set of ablation studies to examine the contribution of several key components in NOLO. More specifically, we aim to answer the following questions.
Does the context video provide informative cues for navigation? Yes. We conduct an ablation study denoted as NOLO-C, in which we evaluate NOLO without context videos in RoboTHOR and Habitat with other configurations unchanged. From Tab. 2, a mean performance drop of NOLO-C is perceived both in RoboTHOR and Habitat compared with NOLO due to lack of context information.
Does temporal coherence loss contribute to better visual representation and performance? Yes. We examine the effect of the temporal coherence loss in Eq. 5. An ablation variant, named NOLO-T, is conducted by removing the temporal coherence loss. The results are shown in Tab. 2. The performance drops slightly in RoboTHOR while more obviously in Habitat. To figure out the underlying reasons for their discrepancy in performance, we further compare the learned visual representation of NOLO and NOLO-T. We randomly select a context video in RoboTHOR and utilize t-SNE projection to visualize their embedding sequences encoded by $E_{\zeta}(f_{t})$ in Fig. 8 in Appendix 8. The embeddings learned by NOLO-T exhibit more irregular stochasticity. In contrast, NOLO learns a better-structured embedding, which demonstrates the effectiveness of the temporal coherence loss. See Appendix 9 for details.
Can NOLO incorporate other off-the-shelf modules to decode actions from videos? Yes. We conduct an ablation, entitled NOLO(M), to decode actions from SuperGlue [70] via point matching instead of GMFlow [89]. We observe an accuracy drop in action decoding (from 92.44% to 80.43%). We then follow the same pipeline and report the performance in Tab. 2. NOLO(M) exhibits around 5% decay in average success rate in RoboTHOR and 7% in Habitat. Therefore, to achieve more satisfactory in-context generalization, we choose GMFlow as the action decoder for NOLO. However, as a general two-stage framework for in-context learning from videos, NOLO is not limited to any kind of pseudo action labeling methods and is capable of incorporating more advanced modules for better performance.
Does NOLO understand the relationship between goal and context video? Yes. We use Grad-CAM [73] saliency maps of the visual encoder to illustrate NOLO’s ability to connect the goal with the context video. As shown in Fig. 6, when a goal image is provided, NOLO accurately identifies key semantic cues in the goal and focuses on related areas in the context frame, indicating its capacity to link goal and context. Without the goal image, however, NOLO lacks this focus, underscoring its reliance on goal-driven context understanding. This result demonstrates NOLO’s ability to interpret the goal and capture relevant information in the context video.
4.6 Real-World Experiments
We constructed a real-world maze environment, as shown in Fig. 7. We employed the Unitree Go2 robot, equipped with a RealSense D435i camera to capture RGB observations. In the maze environment, the robot’s actions are discretized into moving forward by 15 cm or turning 30 degrees at each step. Six target objects–a book, box, cola can, cup, glue, and novel–are placed in the maze. Images of each target are taken from random nearby perspectives. A fixed, non-exploratory, rule-based policy was employed to collect two 600-step video clips in the maze, one for training and the other as context video during policy deployment. For evaluation, each episode consists of 100 steps, with a success radius set at 50 cm. Fig. 5 illustrates key frames from successful trajectories during in-context policy deployment. The robot demonstrated impressive capabilities in navigating straight passages, executing turns at corners, and ultimately reaching the goals by comprehending the dynamics and topological structure inferred from the unseen context video and its relationship to the goal images.
5 Conclusion and Limitaiton
Figure 6: Comparision of saliency maps between NOLO and NOLO-C about context frames with respect to goal frames in RoboTHOR and Habitat.
Figure 7: Setups of real-world maze. The robotic dog mounted with a head camera acts as the navigation agent and six annotated objects are goals.
[…]
Supplementary Material
6 Methodology Details
6.1 Pseudo Action Labeling
We deploy optical flow and filter dominant vectors with the help of the flow map. Theoretically, discretizing filtered vectors into several action categories can be achieved by any unsupervised pattern recognition methods, such as clustering. In practice, we use a simple threshold-based heuristic. We observe that the feature of moving forward is less prominent than turning left or right. Consequently, we focus more on the horizontal displacement of the pixels in the flow map. If the horizontal component of horizontal displacement exceeds a threshold $\tau_{x}$ and the vertical component stays within a threshold $\tau_{y}$ then we classify it as a turning action, or it is considered as a forwarding action.
Occasionally, we notice that the displacement of pixels observed across frames manifests irregular patterns. To mitigate this issue, for each pair of adjacent frames $(f_{t},f_{t+1})$, we filter dominant vectors $\nu_{t}$ that exhibit gradient amplitudes within the upper decile, which are subsequently discretized into actions derived from the action space, making it possible to trace the trajectory of the agent over time.
6.2 Policy Inference
We choose the threshold of choosing action $\beta=0.5$After the loss of BCQ converges, we make decisions based on the learned policy and Q-values. During inference, an action is chosen based on the Q-values weighted by masked action distribution $\pi_{\theta}\left(a_{t}|g,o_{t},\mathcal{T}^{i}\right)$ with probability $\varepsilon=0.999$ and otherwise randomly. Besides, as long as $P_{D}$ outputs a strong signal for termination, the agent will directly output action $STOP$.
6.3 Semantic Actions
[…]
6.4 Model Architecture
[…]
7 Experiment Details
7.1 Dataset Division
[…]
Table 3: Dataset division of RoboTHOR and Habitat
Train
S9hNv5qa7GM XcA2TqTSSAj PX4nDJXEHrG
5LpN3gDmAk7 WYY7iVyf5p8 5q7pvUzZiYa
VFuaQ6m2Qom EDJbREhghzL rPc6DW4iMge
2n8kARJN3HM HxpKQynjfin PuKPg4mmafe
759xd9YjKW5 aayBHfsNo7d jtcxE69GiFV
B6ByNegPMKs p5wJjkQkbXX D7N2EKCX4Sj
1LXtFkjw3qL YVUC4YcDtcY YmJkqBEsHnH
7y3sRwLe3Va r47D5H71a5s vyrNrziPKCB
gxdoqLR6rwA 1pXnuDYAj8r VLzqgDo317F
kEZ7cmS4wCh sT4fr6TAbpF ARNzJeq3xxb
Vvot9Ly1tCj cV4RVeZvu5T VVfe2KiqLaN
jh4fc5c5qoQ 2t7WUuJeko7 D7G3Y4RVNrH
uNb9QFRL6hY V2XKFyX4ASd e9zR4mvMWw7
RPmz2sHmrrY E9uDoFAP3SH 17DRP5sb8fy
JeFG25nYj2p mJXqzFtmKg4 YFuZgdQ5vWj
ac26ZMwG7aT JF19kD82Mey Pm6F8kyY3z2
sKLMLpTHeUy wc2JMjhGNzB dhjEzFoUFzH
82sE5b5pLXE gZ6f7yhEvPG 8WUmhLawc2A
5ZKStnWn8Zo pRbA3pwrgk9 29hnd4uzFmX
GdvgFV5R1Z5 ULsKaCPVFJR q9vSo1VnCiC
Uxmj2M2itWa qoiz87JEwZ2 s8pcmisQ38h
ur6pFq6Qu1A fzynW3qQPVF ZMojNkEp431
r1Q1Z4BcV1o b8cTxDM8gDG i5noydFURQK
JmbYfDe2QKZ
Eval
2azQ1b91cZZ x8F5xyUWy9e QUCTc6BB5sX
Vt2qJdWjCF2 gTV8FGcVJC9 EU6Fwq7SyZv
zsNo4HB9uLZ X7HyMhZNoso yqstnuAEVhm
SN83YJsR3w2 pLe4wQe7qrG 8194nk5LbLH
Z6MFQCViBuw pa4otMbVnkk VzqfbhrpDEA
TbHJrupSAjP oLBMNvg9in8 rqfALeAoiTq
UwV83HsGsw3 gYvKGZ5eRqb
7.2 LMM Baselines
For both GPT-4o and Video-LLaVA, we maintain a queue of size $K=16$ to incorporate sampled context frames, current observation, and goal image into a video clip and designed a consistent prompt to facilitate their understanding of the video content and to guide them to provide a reasonable action at each timestep. The prompt is structured as follows:
These frames pertain to a navigation task. The goal of the video is represented by the last frame, and the second-to-last frame is the current observation. Choose the best action from [move forward, turn left, turn right, stop] that will help achieve the goal depicted in the last frame of the video. Please provide the selected action directly.
We then utilize regular expressions to extract actions from the received language responses. In case of invalid output, we attempt up to three retries, and a random action is selected as output if all attempts fail.
7.3 Hyperparameters
[…]
Hyperparameter Value
Optimizer AdamW
Learning Rate 1e-4
Betas (0.9,0.95)
Weight Decay 0.1
Batch size 26
SA Layers 9
CA Layers 4
Action Embedding Dimension 256
Visual Embedding Dimension 512
Hidden State Dimension 768
BCQ Clip Threshold 0.5
Type of GPUs A100 & Nvidia RTX 4090 Ti
7.4 Computation Resources
For both RoboTHOR and Habitat, we train NOLO and all NOLO-variants on a single 40G A100 GPU for around 1 day. We evaluate NOLO across all created testing tasks (single thread) on a 24G Nvidia RTX 4090 Ti GPU for about 6 hours in RoboTHOR and 5 hours in Habitat.
8 Additional Visualizations
Figure 8: Comparision of context embedding sequences of a sampled video from RoboTHOR between NOLO and NOLO-T.
Sec. 8 compares the t-SNE visual embedding sequence of a context video in RoboTHOR between NOLO and NOLO-T.
9 Evaluation Results
Tables below outline the average performance of different methods in RoboTHOR and Habitat as well as cross evaluation results over 3 random seeds.
9.1 RoboTHOR Evaluation Performance
We also examined the impact of NOLO compared to LMM baselines and visual navigation baselines on two key metrics for vision-based navigation tasks: Success Rate (SR) and Success weighted by normalized inverse Path Length (SPL), as illustrated in the figures in Fig. 9 and Fig. 10. Figs. 9(a), 9(b), 10(a) and 10(b) show the performance in the unseen layout testing set. Figs. 9(c), 9(d), 10(c) and 10(d) show the evaluation results in the unseen room testing set, and The detailed performances of NOLO, random policy, GPT-4o, Video-LLaVA, AVDC, VGM, ZSON, NOLO-C, NOLO-T, NOLO(M) in RoboTHOR are reported in Tabs. 5, 6, 7, 8, 9, 10, 11, 12, 13 and 14 respectively.
(a) Unseen Layout-SR
(b) Unseen Layout-SPL
(c) Unseen Room-SR
(d) Unseen Room-SPL
Figure 9: Compare NOLO with random policy and LMM baselines in RoboTHOR seen and unseen rooms.
(a) Unseen Layout-SR
(b) Unseen Layout-SPL
(c) Unseen Room-SR
(d) Unseen Room-SPL
Figure 10: Compare NOLO with visual navigation baselines in RoboTHOR seen and unseen rooms.
Table 5: NOLO performance on RoboTHOR testing set. “T” denotes FloorPlan_Train and “V” denotes FloorPlan_Val.
[…]
T10_5 77.45$\pm$ 4.36 28.21$\pm$ 1.58 216.35$\pm$ 14.21 1.68$\pm$ 0.13
T11_5 76.44$\pm$ 3.82 36.19$\pm$ 2.32 218.66$\pm$ 13.06 1.62$\pm$ 0.16
T12_5 76.27$\pm$ 3.54 31.89$\pm$ 0.94 228.36$\pm$ 4.96 1.50$\pm$ 0.09
T1_5 59.72$\pm$ 2.39 33.15$\pm$ 1.72 260.69$\pm$ 9.97 2.06$\pm$ 0.05
T2_5 72.54$\pm$ 0.81 28.83$\pm$ 1.45 231.64$\pm$ 4.56 1.76$\pm$ 0.11
T3_5 60.20$\pm$ 1.21 23.29$\pm$ 0.21 289.70$\pm$ 6.56 2.34$\pm$ 0.04
T4_5 77.42$\pm$ 2.74 27.64$\pm$ 1.66 216.64$\pm$ 6.83 1.71$\pm$ 0.13
T5_5 77.58$\pm$ 2.27 32.03$\pm$ 1.45 227.46$\pm$ 4.43 1.54$\pm$ 0.09
T6_5 74.44$\pm$ 4.37 30.13$\pm$ 1.04 217.56$\pm$ 6.36 1.90$\pm$ 0.21
T7_5 80.00$\pm$ 1.77 22.41$\pm$ 1.20 239.99$\pm$ 18.16 1.39$\pm$ 0.11
T8_5 66.25$\pm$ 8.84 27.60$\pm$ 3.87 249.83$\pm$ 31.73 2.07$\pm$ 0.35
T9_5 64.67$\pm$ 3.40 29.72$\pm$ 1.78 270.19$\pm$ 12.79 1.95$\pm$ 0.11
Average 71.92 29.26 238.92 1.79
V1_1 70.83$\pm$ 3.06 30.72$\pm$ 0.78 241.85$\pm$ 11.56 1.74$\pm$ 0.17
V1_2 60.45$\pm$ 3.04 23.57$\pm$ 1.53 285.97$\pm$ 5.54 2.31$\pm$ 0.07
V1_3 82.98$\pm$ 2.37 36.90$\pm$ 2.38 178.16$\pm$ 11.44 1.41$\pm$ 0.11
V1_4 66.85$\pm$ 2.05 26.25$\pm$ 2.42 277.40$\pm$ 15.40 1.85$\pm$ 0.19
V1_5 71.81$\pm$ 2.27 34.42$\pm$ 0.91 238.96$\pm$ 5.10 1.87$\pm$ 0.12
V2_1 56.97$\pm$ 2.61 17.31$\pm$ 0.98 326.74$\pm$ 9.13 2.30$\pm$ 0.20
V2_2 70.56$\pm$ 5.93 31.03$\pm$ 1.34 241.85$\pm$ 24.72 1.92$\pm$ 0.17
V2_3 76.41$\pm$ 3.16 33.33$\pm$ 2.68 216.07$\pm$ 8.19 1.49$\pm$ 0.11
V2_4 64.36$\pm$ 3.57 25.24$\pm$ 0.84 263.61$\pm$ 15.84 2.04$\pm$ 0.15
V2_5 63.33$\pm$ 3.02 24.99$\pm$ 0.36 284.43$\pm$ 5.26 2.44$\pm$ 0.12
V3_1 72.86$\pm$ 1.03 26.81$\pm$ 2.75 240.61$\pm$ 8.40 1.90$\pm$ 0.07
V3_2 79.31$\pm$ 4.62 27.97$\pm$ 1.41 213.71$\pm$ 14.15 1.60$\pm$ 0.23
V3_3 81.48$\pm$ 3.02 30.05$\pm$ 1.48 213.88$\pm$ 10.77 1.46$\pm$ 0.08
V3_4 63.81$\pm$ 5.10 18.56$\pm$ 0.80 283.22$\pm$ 10.86 2.11$\pm$ 0.27
V3_5 75.14$\pm$ 0.20 28.88$\pm$ 2.25 220.18$\pm$ 9.62 1.65$\pm$ 0.04
Average 70.48 27.74 248.44 1.87
Table 6: Random performance on RoboTHOR testing set. “T” denotes FloorPlan_Train and “V” denotes FloorPlan_Val.
[…]
T10_5 27.15$\pm$ 0.50 16.55$\pm$ 0.85 371.96$\pm$ 3.07 2.84$\pm$ 0.00
T11_5 22.83$\pm$ 4.50 11.98$\pm$ 2.67 406.66$\pm$ 15.35 3.24$\pm$ 0.22
T12_5 20.68$\pm$ 1.26 11.93$\pm$ 0.51 413.58$\pm$ 4.25 3.02$\pm$ 0.04
T1_5 22.00$\pm$ 2.00 13.41$\pm$ 1.47 396.63$\pm$ 0.20 3.08$\pm$ 0.09
T2_5 32.83$\pm$ 0.93 14.00$\pm$ 0.57 369.18$\pm$ 2.40 2.54$\pm$ 0.06
T3_5 21.26$\pm$ 6.38 10.25$\pm$ 0.81 421.76$\pm$ 1.69 3.36$\pm$ 0.09
T4_5 21.32$\pm$ 2.23 9.23$\pm$ 1.62 415.40$\pm$ 9.71 3.08$\pm$ 0.06
T5_5 34.95$\pm$ 4.14 16.82$\pm$ 1.57 373.65$\pm$ 12.36 2.55$\pm$ 0.16
T6_5 33.67$\pm$ 2.00 17.65$\pm$ 1.43 357.99$\pm$ 0.93 2.76$\pm$ 0.05
T7_5 30.75$\pm$ 0.75 12.89$\pm$ 0.42 391.34$\pm$ 6.28 2.54$\pm$ 0.12
T8_5 27.00$\pm$ 1.75 11.85$\pm$ 1.68 407.04$\pm$ 8.17 3.16$\pm$ 0.17
T9_5 17.50$\pm$ 1.50 12.73$\pm$ 0.25 412.05$\pm$ 6.07 2.96$\pm$ 0.02
Average 26.00 13.27 394.77 2.93
V1_1 14.50$\pm$ 2.50 9.64$\pm$ 0.75 436.59$\pm$ 3.65 3.13$\pm$ 0.22
V1_2 14.50$\pm$ 4.50 8.63$\pm$ 1.24 433.62$\pm$ 8.48 3.25$\pm$ 0.32
V1_3 31.61$\pm$ 6.87 16.30$\pm$ 0.77 382.54$\pm$ 14.30 2.77$\pm$ 0.13
V1_4 12.28$\pm$ 3.28 8.49$\pm$ 0.73 436.72$\pm$ 9.52 3.18$\pm$ 0.11
V1_5 16.17$\pm$ 1.33 10.29$\pm$ 0.50 424.95$\pm$ 9.29 3.23$\pm$ 0.12
V2_1 5.41$\pm$ 3.68 4.71$\pm$ 1.47 464.95$\pm$ 11.81 3.17$\pm$ 0.03
V2_2 25.61$\pm$ 0.06 13.89$\pm$ 0.33 395.96$\pm$ 1.47 3.17$\pm$ 0.12
V2_3 31.04$\pm$ 2.04 17.05$\pm$ 0.02 377.52$\pm$ 5.43 2.17$\pm$ 0.06
V2_4 23.73$\pm$ 3.96 14.14$\pm$ 0.76 398.90$\pm$ 3.78 2.50$\pm$ 0.10
V2_5 18.20$\pm$ 1.37 11.74$\pm$ 0.15 417.68$\pm$ 0.58 3.25$\pm$ 0.08
V3_1 30.45$\pm$ 0.50 13.00$\pm$ 0.84 381.07$\pm$ 4.86 2.55$\pm$ 0.01
V3_2 34.08$\pm$ 1.75 12.83$\pm$ 0.82 368.68$\pm$ 1.57 2.79$\pm$ 0.06
V3_3 28.39$\pm$ 0.87 11.83$\pm$ 0.51 394.93$\pm$ 4.61 2.84$\pm$ 0.05
V3_4 25.21$\pm$ 3.83 10.30$\pm$ 1.26 397.19$\pm$ 9.02 3.01$\pm$ 0.15
V3_5 35.33$\pm$ 3.00 17.42$\pm$ 1.22 376.74$\pm$ 3.21 2.34$\pm$ 0.07
Average 23.10 12.02 405.87 2.89
Table 7: GPT-4o performance on RoboTHOR testing set. “T” denotes FloorPlan_Train and “V” denotes FloorPlan_Val.
[…]
T10_5 35.29 19.09 358.81 3.00
T11_5 36.00 18.32 363.55 3.15
T12_5 35.29 20.42 359.38 3.17
T1_5 33.33 21.14 362.02 2.78
T2_5 28.57 13.10 367.66 2.76
T3_5 20.00 10.86 377.68 3.30
T4_5 37.27 18.58 357.89 3.12
T5_5 32.73 18.66 363.37 2.95
T6_5 41.67 21.92 351.87 3.24
T7_5 21.25 8.14 374.45 3.10
T8_5 36.25 19.38 356.12 3.07
T9_5 26.00 16.57 366.02 2.72
Average 31.97 17.18 363.23 3.03
V1_1 24.00 13.84 374.19 3.45
V1_2 21.82 11.66 373.75 3.22
V1_3 46.32 26.44 342.01 2.53
V1_4 15.56 8.23 385.98 3.48
V1_5 16.59 5.07 390.28 3.10
V2_1 28.43 17.75 365.67 3.31
V2_2 48.95 23.77 346.67 2.09
V2_3 58.52 30.56 329.43 1.82
V2_4 35.23 20.31 360.92 2.09
V2_5 26.41 14.60 370.46 3.04
V3_1 7.48 2.94 391.37 4.18
V3_2 41.67 21.22 349.08 2.99
V3_3 31.11 13.08 369.33 2.73
V3_4 25.71 11.31 373.28 3.22
V3_5 47.50 26.64 346.99 2.09
Average 31.69 16.49 364.63 2.89
Table 8: Video-LLaVA performance on RoboTHOR testing set. “T” denotes FloorPlan_Train and “V” denotes FloorPlan_Val.
[…]
T10_5 25.88 13.74 370.67 2.86
T11_5 22.67 8.79 380.99 3.58
T12_5 17.65 8.67 379.87 3.30
T1_5 20.00 7.78 383.98 2.77
T2_5 20.95 8.18 382.31 2.67
T3_5 14.12 5.89 388.76 3.84
T4_5 19.09 6.39 385.62 3.31
T5_5 18.18 7.55 385.68 3.04
T6_5 23.84 8.69 378.77 2.78
T7_5 19.86 7.67 382.04 3.01
T8_5 15.79 7.32 384.94 2.84
T9_5 22.83 9.11 379.34 2.98
Average 20.07 8.31 381.92 3.08
V1_1 25.92 10.42 374.65 3.01
V1_2 13.64 5.17 389.82 3.44
V1_3 10.92 1.66 396.36 3.65
V1_4 20.47 8.40 382.13 3.13
V1_5 9.17 3.28 395.26 3.63
V2_1 14.21 6.74 388.84 3.61
V2_2 22.12 8.90 377.86 3.30
V2_3 26.38 12.39 372.44 3.13
V2_4 13.53 4.34 391.22 4.01
V2_5 20.85 6.12 387.35 3.21
V3_1 20.00 6.59 381.67 2.93
V3_2 17.46 6.13 387.79 2.94
V3_3 16.29 5.24 388.70 2.91
V3_4 14.36 5.76 388.94 3.55
V3_5 17.38 7.51 382.77 3.34
Average 17.51 6.58 385.72 3.32
Table 9: AVDC performance on RoboTHOR testing set. T” denotes FloorPlan_Train and V” denotes FloorPlan_Val.
[…]
T10_5 33.84 9.29 422.72 1.92
T11_5 36.72 11.13 434.58 2.22
T12_5 31.24 8.14 437.61 2.12
T1_5 28.98 7.44 428.73 2.00
T2_5 32.59 9.05 440.12 2.14
T3_5 24.53 8.26 444.36 2.45
T4_5 35.21 10.57 423.94 1.99
T5_5 33.47 10.36 433.51 2.09
T6_5 39.12 11.98 415.89 2.32
T7_5 25.74 6.92 439.82 2.18
T8_5 36.23 11.45 421.16 2.27
T9_5 32.67 8.54 426.75 2.15
Average 32.53 9.43 430.77 2.15
V1_1 27.14 8.21 462.48 2.65
V1_2 29.62 11.18 440.87 2.53
V1_3 30.28 15.24 405.27 1.84
V1_4 22.84 7.91 454.28 2.59
V1_5 30.41 6.76 447.91 2.21
V2_1 34.32 13.43 425.68 2.05
V2_2 48.45 10.92 398.67 1.25
V2_3 43.67 12.57 387.45 1.73
V2_4 36.18 14.97 404.22 1.49
V2_5 31.84 12.98 412.88 2.29
V3_1 11.14 5.23 450.75 3.07
V3_2 28.89 19.34 409.38 1.97
V3_3 34.67 13.15 429.52 2.21
V3_4 27.93 10.27 439.74 2.55
V3_5 29.78 11.33 399.28 1.53
Average 31.14 11.57 424.56 2.13
Table 10: VGM performance on RoboTHOR testing set. T” denotes FloorPlan_Train and V” denotes FloorPlan_Val.
[…]
T10_5 46.02 25.47 375.88 2.49
T11_5 44.98 23.65 392.45 2.71
T12_5 50.13 22.71 378.29 2.39
T1_5 49.75 26.02 371.15 2.55
T2_5 43.59 21.54 388.67 2.23
T3_5 46.93 22.88 367.36 2.67
T4_5 48.32 24.07 381.74 2.41
T5_5 44.25 24.61 373.28 2.44
T6_5 50.41 25.19 369.24 2.33
T7_5 45.67 22.32 365.72 2.62
T8_5 46.89 23.78 380.91 2.35
T9_5 47.54 24.56 378.32 2.13
Average 47.04 23.90 376.92 2.44
V1_1 41.13 21.57 405.87 2.64
V1_2 39.23 19.31 412.88 2.87
V1_3 43.67 22.13 386.24 1.77
V1_4 42.45 18.27 416.54 2.41
V1_5 35.88 18.68 401.96 2.54
V2_1 44.26 20.67 382.33 2.38
V2_2 49.87 23.12 394.56 1.85
V2_3 41.22 21.87 385.19 1.91
V2_4 38.96 22.07 396.72 2.25
V2_5 43.78 20.52 405.33 2.64
V3_1 35.37 12.43 420.64 3.22
V3_2 39.16 21.38 389.33 2.42
V3_3 33.78 18.04 406.57 2.11
V3_4 30.15 14.92 416.38 2.82
V3_5 48.76 24.87 392.94 1.87
Average 40.51 19.99 400.90 2.38
Table 11: ZSON performance on RoboTHOR testing set. T” denotes FloorPlan_Train and V” denotes FloorPlan_Val.
[…]
T10_5 38.89 20.69 350.67 2.23
T11_5 35.06 19.92 362.44 2.41
T12_5 40.06 22.54 345.23 2.16
T1_5 36.17 19.45 358.91 2.09
T2_5 34.37 22.58 351.12 2.02
T3_5 41.86 23.24 340.11 2.34
T4_5 38.72 21.52 346.72 2.29
T5_5 36.06 20.35 349.33 2.25
T6_5 43.17 22.61 335.84 2.38
T7_5 35.17 20.45 360.17 2.15
T8_5 35.96 20.27 348.52 2.31
T9_5 38.52 22.08 342.66 2.18
Average 37.83 21.31 349.31 2.23
V1_1 37.91 18.12 378.87 2.72
V1_2 34.50 17.64 386.19 2.91
V1_3 43.06 20.88 359.72 1.77
V1_4 28.71 16.29 395.18 2.56
V1_5 33.77 15.17 380.14 2.64
V2_1 40.93 19.42 361.48 2.12
V2_2 41.60 25.01 353.26 1.43
V2_3 43.34 21.41 368.11 1.85
V2_4 39.55 22.75 354.24 2.25
V2_5 40.02 19.83 360.73 2.72
V3_1 19.52 10.81 392.54 3.22
V3_2 37.89 19.64 368.18 2.56
V3_3 32.41 15.28 382.45 2.21
V3_4 23.77 13.04 390.76 2.89
V3_5 39.79 26.81 358.32 1.73
Average 35.78 18.81 372.68 2.37
Table 12: NOLO-C performance on RoboTHOR testing set. “T” denotes FloorPlan_Train and “V” denotes FloorPlan_Val.
[…]
T10_5 72.94$\pm$ 5.08 27.83$\pm$ 0.15 252.83$\pm$ 8.62 1.86$\pm$ 0.29
T11_5 82.00$\pm$ 1.44 36.28$\pm$ 2.90 195.77$\pm$ 8.38 1.50$\pm$ 0.05
T12_5 73.53$\pm$ 0.83 32.13$\pm$ 1.26 230.53$\pm$ 8.63 1.59$\pm$ 0.05
T1_5 64.72$\pm$ 1.04 32.41$\pm$ 1.39 258.61$\pm$ 2.70 1.94$\pm$ 0.06
T2_5 70.63$\pm$ 3.40 29.21$\pm$ 3.18 243.38$\pm$ 16.56 1.78$\pm$ 0.11
T3_5 62.16$\pm$ 1.94 23.55$\pm$ 0.62 288.69$\pm$ 8.45 2.27$\pm$ 0.09
T4_5 72.12$\pm$ 5.77 27.65$\pm$ 0.21 244.76$\pm$ 16.37 1.89$\pm$ 0.20
T5_5 75.61$\pm$ 0.57 32.77$\pm$ 1.68 232.83$\pm$ 3.66 1.60$\pm$ 0.04
T6_5 67.78$\pm$ 4.37 27.60$\pm$ 0.31 254.15$\pm$ 5.21 2.00$\pm$ 0.12
T7_5 75.62$\pm$ 1.84 19.90$\pm$ 2.02 256.17$\pm$ 8.42 1.55$\pm$ 0.03
T8_5 65.62$\pm$ 2.55 26.02$\pm$ 2.55 253.82$\pm$ 8.94 2.00$\pm$ 0.07
T9_5 56.67$\pm$ 4.11 24.40$\pm$ 2.52 303.16$\pm$ 8.15 2.18$\pm$ 0.05
Average 69.95 28.31 251.23 1.85
V1_1 64.33$\pm$ 5.44 25.29$\pm$ 5.10 283.93$\pm$ 29.92 2.12$\pm$ 0.22
V1_2 57.42$\pm$ 3.57 22.97$\pm$ 1.36 305.57$\pm$ 10.25 2.34$\pm$ 0.15
V1_3 82.98$\pm$ 1.38 41.56$\pm$ 1.40 197.25$\pm$ 10.54 1.43$\pm$ 0.04
V1_4 62.22$\pm$ 3.27 25.73$\pm$ 1.58 284.70$\pm$ 10.51 2.26$\pm$ 0.10
V1_5 58.61$\pm$ 1.42 31.95$\pm$ 2.50 276.70$\pm$ 4.67 2.53$\pm$ 0.12
V2_1 52.73$\pm$ 1.96 17.41$\pm$ 1.69 336.15$\pm$ 1.47 2.59$\pm$ 0.08
V2_2 67.22$\pm$ 4.33 28.52$\pm$ 2.17 268.21$\pm$ 2.39 1.95$\pm$ 0.07
V2_3 79.74$\pm$ 3.46 35.81$\pm$ 0.85 209.43$\pm$ 23.21 1.34$\pm$ 0.08
V2_4 62.95$\pm$ 6.01 25.65$\pm$ 0.77 276.85$\pm$ 16.89 2.10$\pm$ 0.19
V2_5 58.70$\pm$ 0.94 24.36$\pm$ 1.83 290.25$\pm$ 2.98 2.55$\pm$ 0.05
V3_1 71.43$\pm$ 2.33 26.32$\pm$ 3.77 240.97$\pm$ 20.02 2.00$\pm$ 0.14
V3_2 71.53$\pm$ 1.75 26.78$\pm$ 0.32 238.58$\pm$ 3.53 1.99$\pm$ 0.09
V3_3 72.59$\pm$ 4.57 27.82$\pm$ 1.74 242.52$\pm$ 9.40 1.73$\pm$ 0.18
V3_4 57.30$\pm$ 2.65 18.35$\pm$ 0.51 300.88$\pm$ 11.16 2.30$\pm$ 0.08
V3_5 70.97$\pm$ 3.16 29.52$\pm$ 1.81 243.77$\pm$ 19.50 1.96$\pm$ 0.14
Average 66.05 27.20 266.38 2.08
Table 13: NOLO-T performance on RoboTHOR testing set. “T” denotes FloorPlan_Train and “V” denotes FloorPlan_Val.
[…]
T10_5 76.47$\pm$ 3.53 29.97$\pm$ 2.31 225.76$\pm$ 18.05 1.71$\pm$ 0.16
T11_5 78.00$\pm$ 0.67 36.00$\pm$ 0.43 210.91$\pm$ 7.94 1.56$\pm$ 0.10
T12_5 75.88$\pm$ 0.59 30.77$\pm$ 1.29 233.25$\pm$ 4.82 1.62$\pm$ 0.02
T1_5 64.17$\pm$ 2.50 32.47$\pm$ 4.30 254.97$\pm$ 6.37 2.07$\pm$ 0.01
T2_5 65.71$\pm$ 6.67 26.12$\pm$ 4.84 259.38$\pm$ 29.69 1.95$\pm$ 0.25
T3_5 59.41$\pm$ 4.12 22.67$\pm$ 0.98 303.37$\pm$ 11.79 2.39$\pm$ 0.08
T4_5 72.73$\pm$ 0.91 27.02$\pm$ 2.10 238.05$\pm$ 10.76 1.89$\pm$ 0.09
T5_5 78.18$\pm$ 3.64 33.88$\pm$ 0.94 224.37$\pm$ 22.06 1.48$\pm$ 0.12
T6_5 74.17$\pm$ 0.83 32.28$\pm$ 1.91 216.92$\pm$ 17.75 1.83$\pm$ 0.11
T7_5 80.62$\pm$ 0.62 23.85$\pm$ 1.03 232.86$\pm$ 15.75 1.38$\pm$ 0.08
T8_5 66.25$\pm$ 2.50 30.95$\pm$ 2.62 245.14$\pm$ 15.49 2.06$\pm$ 0.12
T9_5 61.00$\pm$ 1.00 28.92$\pm$ 1.60 285.27$\pm$ 2.75 1.80$\pm$ 0.20
Average 71.05 29.57 244.19 1.81
V1_1 64.00$\pm$ 2.00 26.54$\pm$ 1.03 277.16$\pm$ 12.08 2.25$\pm$ 0.02
V1_2 64.55 25.30$\pm$ 0.81 272.34$\pm$ 12.70 2.20$\pm$ 0.01
V1_3 80.53$\pm$ 5.79 36.93$\pm$ 0.98 201.32$\pm$ 9.73 1.62$\pm$ 0.22
V1_4 59.44$\pm$ 7.22 25.50$\pm$ 1.25 283.97$\pm$ 24.48 2.14$\pm$ 0.21
V1_5 62.50$\pm$ 0.83 32.49$\pm$ 0.23 268.84$\pm$ 5.32 2.06$\pm$ 0.01
V2_1 56.36$\pm$ 0.91 17.59$\pm$ 0.30 329.25$\pm$ 10.75 2.44$\pm$ 0.04
V2_2 65.56$\pm$ 2.22 30.85$\pm$ 2.09 256.13$\pm$ 9.46 2.07$\pm$ 0.07
V2_3 76.92$\pm$ 1.54 34.22$\pm$ 0.03 209.07$\pm$ 4.52 1.59$\pm$ 0.02
V2_4 68.08$\pm$ 0.38 25.24$\pm$ 0.19 265.69$\pm$ 6.21 1.93$\pm$ 0.00
V2_5 65.22$\pm$ 3.48 24.79$\pm$ 0.74 290.28$\pm$ 11.04 2.29$\pm$ 0.14
V3_1 69.52$\pm$ 0.95 25.09$\pm$ 0.17 244.60$\pm$ 10.47 2.04$\pm$ 0.03
V3_2 75.83$\pm$ 0.83 27.14$\pm$ 2.16 223.59$\pm$ 6.15 1.81$\pm$ 0.06
V3_3 73.70$\pm$ 2.59 24.63$\pm$ 2.77 245.72$\pm$ 15.48 1.75$\pm$ 0.15
V3_4 52.38$\pm$ 3.81 17.61$\pm$ 0.82 316.57$\pm$ 4.58 2.46$\pm$ 0.12
V3_5 76.67$\pm$ 0.83 31.29$\pm$ 1.91 211.91$\pm$ 5.29 1.62$\pm$ 0.04
Average 67.42 27.01 259.76 2.02
Table 14: NOLO(M) performance on RoboTHOR testing set. “T” denotes FloorPlan_Train and “V” denotes FloorPlan_Val.
[…]
T10_5 80.59$\pm$ 4.12 32.92$\pm$ 1.93 202.68$\pm$ 22.39 1.47$\pm$ 0.12
T11_5 77.33$\pm$ 2.67 36.00$\pm$ 3.27 218.04$\pm$ 1.64 1.64$\pm$ 0.01
T12_5 78.82 36.62$\pm$ 0.82 211.44$\pm$ 9.89 1.55$\pm$ 0.01
T1_5 50.83$\pm$ 5.83 27.99$\pm$ 2.85 302.14$\pm$ 14.69 2.35$\pm$ 0.10
T2_5 69.05$\pm$ 1.43 25.09$\pm$ 2.83 256.49$\pm$ 1.55 1.85$\pm$ 0.13
T3_5 54.71$\pm$ 0.59 21.49$\pm$ 0.94 312.69$\pm$ 3.18 2.61$\pm$ 0.08
T4_5 73.64$\pm$ 2.73 25.06$\pm$ 1.03 245.16$\pm$ 17.39 1.78$\pm$ 0.08
T5_5 75.45$\pm$ 0.91 31.23$\pm$ 1.14 234.16$\pm$ 1.94 1.75$\pm$ 0.07
T6_5 65.83$\pm$ 2.50 24.81$\pm$ 0.36 276.04$\pm$ 19.72 1.95$\pm$ 0.01
T7_5 70.62$\pm$ 0.63 19.64$\pm$ 0.81 281.29$\pm$ 2.88 1.77$\pm$ 0.05
T8_5 59.38$\pm$ 6.87 23.84$\pm$ 2.27 287.47$\pm$ 17.78 2.38$\pm$ 0.37
T9_5 59.00$\pm$ 7.00 31.86$\pm$ 5.95 278.41$\pm$ 28.49 2.12$\pm$ 0.22
Average 67.94 28.05 258.83 1.93
V1_1 60.00$\pm$ 3.00 25.32$\pm$ 3.23 294.85$\pm$ 21.57 2.29$\pm$ 0.09
V1_2 60.00$\pm$ 1.82 24.32$\pm$ 3.48 291.72$\pm$ 6.24 2.18$\pm$ 0.09
V1_3 82.63$\pm$ 1.58 35.75$\pm$ 3.13 200.15$\pm$ 3.24 1.41$\pm$ 0.09
V1_4 62.22$\pm$ 1.11 24.95$\pm$ 0.16 288.34$\pm$ 6.13 2.19$\pm$ 0.21
V1_5 64.58$\pm$ 4.58 31.77$\pm$ 3.05 265.98$\pm$ 16.41 2.04$\pm$ 0.16
V2_1 49.09$\pm$ 1.82 14.51$\pm$ 2.21 358.64$\pm$ 2.52 2.68$\pm$ 0.06
V2_2 66.11$\pm$ 2.78 30.06$\pm$ 1.07 261.83$\pm$ 10.29 2.11$\pm$ 0.15
V2_3 77.69$\pm$ 0.77 31.33$\pm$ 1.61 232.93$\pm$ 6.27 1.49$\pm$ 0.02
V2_4 57.69$\pm$ 1.54 23.61$\pm$ 0.69 296.56$\pm$ 8.78 2.26$\pm$ 0.07
V2_5 59.57$\pm$ 4.78 23.29$\pm$ 0.57 303.90$\pm$ 19.02 2.48$\pm$ 0.30
V3_1 70.48$\pm$ 1.90 25.14$\pm$ 1.24 254.00$\pm$ 0.10 1.87$\pm$ 0.06
V3_2 71.25$\pm$ 0.42 25.59$\pm$ 0.67 244.68$\pm$ 2.23 1.88$\pm$ 0.04
V3_3 67.78$\pm$ 0.37 24.20$\pm$ 1.78 258.00$\pm$ 2.40 1.89$\pm$ 0.03
V3_4 54.29$\pm$ 0.95 13.15$\pm$ 1.33 330.42$\pm$ 8.02 2.40$\pm$ 0.00
V3_5 67.50$\pm$ 1.67 30.21$\pm$ 4.40 246.60$\pm$ 8.73 2.05$\pm$ 0.04
Average 64.73 25.55 275.24 2.08
9.2 Habitat Evaluation Performance
Also, we examined the impact of NOLO compared to LMM baselines and visual navigation baselines on two key metrics for vision-based navigation tasks: Success Rate (SR) and Success weighted by normalized inverse Path Length (SPL), as illustrated in Fig. 11. The performances of NOLO, random policy, GPT-4o, Video-LLaVA, AVDC, VGM, ZSON, NOLO-C, NOLO-T, NOLO(M) in Habitat testing scenes are reported in Tabs. 15, 16, 17, 18, 19, 20, 21, 22, 23 and 24 respectively.
(a) Average SR
(b) Average SPL
(c) Average SR
(d) Average SPL
Figure 11: Compare NOLO with random policy, LMM baselines and visual navigation baselines in 20 Habitat testing scenes.
Table 15: NOLO performance in Habitat testing scenes.
[…]
2azQ1b91cZZ 50.26$\pm$ 3.10 28.99$\pm$ 1.40 311.44$\pm$ 11.24 2.80$\pm$ 0.12
8194nk5LbLH 43.94$\pm$ 1.55 22.60$\pm$ 0.51 341.28$\pm$ 1.28 3.10$\pm$ 0.08
EU6Fwq7SyZv 42.33$\pm$ 1.70 18.15$\pm$ 1.13 362.99$\pm$ 5.59 2.90$\pm$ 0.06
QUCTc6BB5sX 44.67$\pm$ 6.24 21.94$\pm$ 1.36 344.75$\pm$ 21.03 4.25$\pm$ 0.84
SN83YJsR3w2 38.33$\pm$ 6.24 15.74$\pm$ 2.48 379.77$\pm$ 17.81 2.97$\pm$ 0.97
TbHJrupSAjP 36.67$\pm$ 10.19 14.29$\pm$ 3.23 376.57$\pm$ 28.90 3.85$\pm$ 0.79
UwV83HsGsw3 30.48$\pm$ 3.75 13.66$\pm$ 0.67 402.48$\pm$ 14.42 4.80$\pm$ 0.58
Vt2qJdWjCF2 34.58$\pm$ 4.60 12.75$\pm$ 4.26 387.33$\pm$ 21.56 6.63$\pm$ 0.81
VzqfbhrpDEA 33.70$\pm$ 1.89 17.94$\pm$ 1.15 377.91$\pm$ 7.07 5.25$\pm$ 0.22
X7HyMhZNoso 66.67$\pm$ 4.38 30.84$\pm$ 2.47 275.02$\pm$ 20.74 1.37$\pm$ 0.22
Z6MFQCViBuw 40.67$\pm$ 6.18 16.25$\pm$ 3.41 369.67$\pm$ 21.35 4.62$\pm$ 0.64
gTV8FGcVJC9 53.67$\pm$ 4.03 29.49$\pm$ 2.20 292.08$\pm$ 12.03 3.98$\pm$ 0.12
gYvKGZ5eRqb 25.00$\pm$ 8.16 7.34$\pm$ 4.04 433.63$\pm$ 30.22 6.06$\pm$ 0.59
oLBMNvg9in8 45.67$\pm$ 4.03 24.45$\pm$ 3.33 335.14$\pm$ 14.15 2.19$\pm$ 0.16
pLe4wQe7qrG 70.00$\pm$ 2.72 38.63$\pm$ 2.66 218.44$\pm$ 6.09 0.88$\pm$ 0.24
pa4otMbVnkk 41.11$\pm$ 11.00 20.99$\pm$ 6.68 346.89$\pm$ 36.77 5.95$\pm$ 1.17
rqfALeAoiTq 47.50$\pm$ 3.12 21.92$\pm$ 2.23 338.93$\pm$ 11.95 2.57$\pm$ 0.11
x8F5xyUWy9e 51.82$\pm$ 0.74 23.36$\pm$ 0.97 320.58$\pm$ 4.80 2.91$\pm$ 0.06
yqstnuAEVhm 38.89$\pm$ 0.79 17.68$\pm$ 1.86 369.21$\pm$ 3.18 2.54$\pm$ 0.12
zsNo4HB9uLZ 37.08$\pm$ 3.28 18.48$\pm$ 2.66 367.31$\pm$ 11.94 3.77$\pm$ 0.34
Average 43.65 20.77 347.57 3.67
Table 16: Random performance in Habitat testing scenes.
[…]
2azQ1b91cZZ 28.73$\pm$ 6.04 13.11$\pm$ 3.30 386.34$\pm$ 24.39 3.91$\pm$ 0.26
8194nk5LbLH 32.68$\pm$ 4.95 11.77$\pm$ 0.59 380.37$\pm$ 10.02 3.78$\pm$ 0.13
EU6Fwq7SyZv 21.50$\pm$ 2.50 8.57$\pm$ 0.97 409.05$\pm$ 12.59 5.07$\pm$ 0.18
QUCTc6BB5sX 21.50$\pm$ 0.50 10.60$\pm$ 1.00 399.77$\pm$ 3.10 5.67$\pm$ 0.00
SN83YJsR3w2 12.00$\pm$ 3.00 5.89$\pm$ 1.05 428.02$\pm$ 12.97 4.97$\pm$ 0.14
TbHJrupSAjP 18.07$\pm$ 0.21 5.57$\pm$ 1.55 427.09$\pm$ 10.29 5.89$\pm$ 0.37
UwV83HsGsw3 15.93$\pm$ 1.93 4.78$\pm$ 1.42 435.46$\pm$ 1.59 5.57$\pm$ 0.16
Vt2qJdWjCF2 20.75$\pm$ 1.75 8.75$\pm$ 0.86 415.93$\pm$ 1.07 6.02$\pm$ 0.05
VzqfbhrpDEA 25.06$\pm$ 2.28 7.90$\pm$ 1.88 412.03$\pm$ 12.59 5.14$\pm$ 0.05
X7HyMhZNoso 45.21$\pm$ 0.21 17.71$\pm$ 0.35 342.76$\pm$ 8.05 2.86$\pm$ 0.03
Z6MFQCViBuw 17.50$\pm$ 4.50 5.40$\pm$ 1.41 430.17$\pm$ 14.77 6.83$\pm$ 0.48
gTV8FGcVJC9 34.00$\pm$ 5.00 13.64$\pm$ 1.80 370.20$\pm$ 16.76 4.76$\pm$ 0.08
gYvKGZ5eRqb 7.00$\pm$ 3.00 5.39$\pm$ 2.65 457.60$\pm$ 20.20 7.21$\pm$ 0.10
oLBMNvg9in8 20.50$\pm$ 0.50 8.57$\pm$ 1.53 418.56$\pm$ 8.32 4.28$\pm$ 0.10
pLe4wQe7qrG 41.17$\pm$ 7.17 15.29$\pm$ 4.11 348.37$\pm$ 24.83 3.17$\pm$ 0.90
pa4otMbVnkk 22.83$\pm$ 7.83 14.06$\pm$ 0.92 391.77$\pm$ 29.37 6.59$\pm$ 0.03
rqfALeAoiTq 22.00$\pm$ 1.33 9.19$\pm$ 2.09 410.70$\pm$ 6.22 4.25$\pm$ 0.36
x8F5xyUWy9e 29.05$\pm$ 3.23 12.57$\pm$ 0.77 388.25$\pm$ 1.57 4.69$\pm$ 0.14
yqstnuAEVhm 23.67$\pm$ 1.17 9.23$\pm$ 0.06 410.58$\pm$ 9.25 3.90$\pm$ 0.13
zsNo4HB9uLZ 31.37$\pm$ 1.13 12.54$\pm$ 0.88 377.44$\pm$ 10.84 4.78$\pm$ 0.14
Average 24.53 10.03 402.02 4.97
Table 17: GPT-4o performance in Habitat testing scenes.
[…]
2azQ1b91cZZ 31.67 17.77 348.79 4.19
8194nk5LbLH 26.67 12.07 360.59 4.21
EU6Fwq7SyZv 37.41 20.31 347.59 5.10
QUCTc6BB5sX 36.90 20.04 347.43 6.38
SN83YJsR3w2 37.24 21.26 344.79 4.69
TbHJrupSAjP 36.89 23.76 339.05 2.99
UwV83HsGsw3 32.86 22.14 344.40 4.97
Vt2qJdWjCF2 32.50 18.00 350.69 7.19
VzqfbhrpDEA 36.67 19.97 346.93 5.26
X7HyMhZNoso 42.86 26.08 335.63 2.66
Z6MFQCViBuw 24.00 12.83 362.88 7.99
gTV8FGcVJC9 38.49 21.29 344.03 4.25
gYvKGZ5eRqb 37.46 23.24 340.84 4.58
oLBMNvg9in8 38.34 23.79 335.41 2.06
pLe4wQe7qrG 29.32 18.54 350.50 5.77
pa4otMbVnkk 34.82 16.11 351.36 5.98
rqfALeAoiTq 39.49 23.56 339.48 3.32
x8F5xyUWy9e 40.91 23.94 338.39 4.23
yqstnuAEVhm 35.00 21.56 345.06 3.41
zsNo4HB9uLZ 33.75 21.45 342.88 4.72
Average 35.16 20.39 345.84 4.70
Table 18: Video-LLaVA performance in Habitat testing scenes.
[…]
2azQ1b91cZZ 20.00 10.21 381.51 3.77
8194nk5LbLH 12.50 7.28 387.78 5.17
EU6Fwq7SyZv 14.00 7.92 385.03 4.04
QUCTc6BB5sX 19.00 12.24 378.96 5.26
SN83YJsR3w2 20.00 12.34 379.00 3.72
TbHJrupSAjP 7.14 3.01 395.36 5.05
UwV83HsGsw3 8.57 4.70 393.87 5.29
Vt2qJdWjCF2 11.32 6.66 388.67 6.24
VzqfbhrpDEA 16.20 10.37 384.72 5.42
X7HyMhZNoso 11.27 4.48 393.52 4.49
Z6MFQCViBuw 24.60 15.43 371.56 4.64
gTV8FGcVJC9 25.00 15.86 371.61 4.50
gYvKGZ5eRqb 10.00 7.19 389.55 5.23
oLBMNvg9in8 14.00 9.90 380.34 4.40
pLe4wQe7qrG 13.33 12.94 377.50 3.17
pa4otMbVnkk 10.00 5.57 387.50 6.02
rqfALeAoiTq 12.50 8.34 386.06 3.91
x8F5xyUWy9e 15.91 11.81 380.44 3.43
yqstnuAEVhm 10.54 6.06 390.36 5.19
zsNo4HB9uLZ 15.01 8.03 387.20 5.30
Average 14.55 9.02 384.53 4.71
Table 19: AVDC performance in Habitat testing scenes.
[…]
2azQ1b91cZZ 24.12 7.93 429.30 3.44
8194nk5LbLH 29.33 14.02 375.06 5.05
EU6Fwq7SyZv 26.95 12.35 442.17 5.45
QUCTc6BB5sX 22.14 10.48 450.81 3.54
SN83YJsR3w2 30.77 15.06 395.85 4.21
TbHJrupSAjP 31.49 10.74 420.21 4.96
UwV83HsGsw3 23.57 12.84 393.14 3.33
Vt2qJdWjCF2 28.05 11.54 426.06 4.42
VzqfbhrpDEA 25.22 14.19 410.55 3.45
X7HyMhZNoso 29.58 12.82 414.27 4.20
Z6MFQCViBuw 26.78 13.78 430.31 4.63
gTV8FGcVJC9 28.29 10.00 411.56 4.27
gYvKGZ5eRqb 26.52 10.18 421.96 4.92
oLBMNvg9in8 25.44 13.00 417.07 3.91
pLe4wQe7qrG 27.88 16.95 429.00 4.40
pa4otMbVnkk 29.31 10.21 434.91 5.12
rqfALeAoiTq 28.45 12.56 403.33 4.42
x8F5xyUWy9e 24.68 9.76 432.66 3.68
yqstnuAEVhm 30.19 13.00 399.91 4.39
zsNo4HB9uLZ 26.94 13.32 424.30 4.14
Average 27.29 12.24 418.12 4.30
Table 20: VGM performance in Habitat testing scenes.
[…]
2azQ1b91cZZ 35.88 16.12 373.11 4.48
8194nk5LbLH 31.27 19.02 423.51 3.85
EU6Fwq7SyZv 34.10 18.74 368.62 4.06
QUCTc6BB5sX 36.27 16.88 367.95 4.29
SN83YJsR3w2 32.92 17.09 426.29 4.73
TbHJrupSAjP 35.08 18.43 395.03 4.40
UwV83HsGsw3 33.45 16.21 424.39 3.82
Vt2qJdWjCF2 36.12 15.91 403.36 4.64
VzqfbhrpDEA 33.57 19.50 397.14 3.67
X7HyMhZNoso 34.88 18.62 398.75 4.05
Z6MFQCViBuw 32.67 15.53 398.10 4.37
gTV8FGcVJC9 35.13 17.68 390.74 4.36
gYvKGZ5eRqb 34.77 16.44 393.34 4.36
oLBMNvg9in8 31.22 17.13 391.16 3.38
pLe4wQe7qrG 36.45 18.04 357.49 4.19
pa4otMbVnkk 34.02 16.66 414.59 3.99
rqfALeAoiTq 35.23 18.21 369.25 4.20
x8F5xyUWy9e 32.45 17.90 380.19 4.32
yqstnuAEVhm 34.99 15.80 412.49 4.17
zsNo4HB9uLZ 36.23 16.09 392.92 4.39
Average 34.34 17.30 393.92 4.19
Table 21: ZSON performance in Habitat testing scenes.
[…]
2azQ1b91cZZ 30.22 21.90 400.05 4.49
8194nk5LbLH 35.14 17.42 397.11 4.13
EU6Fwq7SyZv 31.88 18.15 360.23 4.71
QUCTc6BB5sX 34.62 21.22 379.75 4.42
SN83YJsR3w2 30.87 19.88 415.70 4.21
TbHJrupSAjP 33.94 18.75 428.33 4.05
UwV83HsGsw3 35.56 20.32 391.88 3.92
Vt2qJdWjCF2 33.18 22.58 400.38 4.37
VzqfbhrpDEA 29.34 19.49 393.52 4.72
X7HyMhZNoso 34.90 18.22 376.97 4.29
Z6MFQCViBuw 32.12 21.74 380.55 4.91
gTV8FGcVJC9 30.25 17.96 406.00 4.35
gYvKGZ5eRqb 35.44 22.08 412.43 3.81
oLBMNvg9in8 31.53 19.14 366.82 4.58
pLe4wQe7qrG 30.87 20.69 393.75 4.14
pa4otMbVnkk 33.97 18.56 418.57 4.28
rqfALeAoiTq 34.12 19.24 378.22 4.33
x8F5xyUWy9e 32.78 18.97 416.39 3.91
yqstnuAEVhm 31.41 21.66 419.74 4.27
zsNo4HB9uLZ 34.55 19.31 382.41 4.43
Average 32.83 19.86 395.94 4.32
Table 22: NOLO-C performance in Habitat testing scenes.
[…]
2azQ1b91cZZ 37.18$\pm$ 4.28 22.84$\pm$ 3.63 359.71$\pm$ 17.26 4.38$\pm$ 0.69
8194nk5LbLH 36.97$\pm$ 1.55 19.47$\pm$ 0.99 366.57$\pm$ 4.66 2.97$\pm$ 0.34
EU6Fwq7SyZv 34.00$\pm$ 2.16 18.48$\pm$ 0.86 378.66$\pm$ 2.54 3.18$\pm$ 0.31
QUCTc6BB5sX 40.33$\pm$ 1.25 23.52$\pm$ 0.77 355.46$\pm$ 1.65 4.83$\pm$ 0.32
SN83YJsR3w2 15.00$\pm$ 7.07 6.92$\pm$ 2.91 455.47$\pm$ 15.62 4.56$\pm$ 1.16
TbHJrupSAjP 20.95$\pm$ 5.99 6.05$\pm$ 1.97 437.43$\pm$ 14.11 4.53$\pm$ 0.51
UwV83HsGsw3 26.67$\pm$ 1.35 17.00$\pm$ 1.76 401.74$\pm$ 3.26 4.77$\pm$ 0.41
Vt2qJdWjCF2 19.17$\pm$ 5.62 8.34$\pm$ 0.78 437.86$\pm$ 13.29 11.49$\pm$ 0.85
VzqfbhrpDEA 25.19$\pm$ 1.05 15.50$\pm$ 2.16 398.60$\pm$ 9.64 5.91$\pm$ 1.16
X7HyMhZNoso 59.05$\pm$ 2.94 30.08$\pm$ 0.75 290.51$\pm$ 7.25 2.00$\pm$ 0.27
Z6MFQCViBuw 38.00$\pm$ 5.66 15.43$\pm$ 1.57 375.89$\pm$ 16.03 5.19$\pm$ 0.22
gTV8FGcVJC9 41.33$\pm$ 1.70 25.04$\pm$ 1.93 335.27$\pm$ 13.56 4.63$\pm$ 0.73
gYvKGZ5eRqb 11.67$\pm$ 2.36 3.51$\pm$ 1.45 474.47$\pm$ 10.70 10.91$\pm$ 0.46
oLBMNvg9in8 44.33$\pm$ 2.36 25.58$\pm$ 1.67 332.98$\pm$ 8.66 1.85$\pm$ 0.12
pLe4wQe7qrG 52.22$\pm$ 3.14 26.99$\pm$ 1.25 316.10$\pm$ 10.12 1.54$\pm$ 0.48
pa4otMbVnkk 14.44$\pm$ 5.67 7.67$\pm$ 3.04 447.38$\pm$ 20.68 7.15$\pm$ 1.53
rqfALeAoiTq 36.94$\pm$ 4.63 18.40$\pm$ 1.58 370.27$\pm$ 11.73 2.45$\pm$ 0.07
x8F5xyUWy9e 46.97$\pm$ 4.71 21.66$\pm$ 1.96 335.51$\pm$ 11.35 2.43$\pm$ 0.57
yqstnuAEVhm 38.61$\pm$ 6.17 17.60$\pm$ 3.23 372.49$\pm$ 22.68 2.89$\pm$ 0.55
zsNo4HB9uLZ 32.50$\pm$ 3.54 18.14$\pm$ 0.90 382.85$\pm$ 11.49 3.92$\pm$ 0.54
Average 33.58 17.41 381.26 4.58
Table 23: NOLO-T performance in Habitat testing scenes.
[…]
2azQ1b91cZZ 41.28$\pm$ 2.97 25.50$\pm$ 2.75 344.03$\pm$ 15.14 4.45$\pm$ 0.08
8194nk5LbLH 39.70$\pm$ 2.61 19.67$\pm$ 0.74 364.86$\pm$ 3.70 3.04$\pm$ 0.27
EU6Fwq7SyZv 41.00$\pm$ 3.27 21.33$\pm$ 2.90 357.03$\pm$ 11.76 3.37$\pm$ 0.17
QUCTc6BB5sX 41.50$\pm$ 4.50 21.52$\pm$ 2.39 357.77$\pm$ 11.40 5.51$\pm$ 0.73
SN83YJsR3w2 21.67$\pm$ 11.79 7.83$\pm$ 4.20 432.92$\pm$ 32.98 5.79$\pm$ 1.89
TbHJrupSAjP 29.52$\pm$ 0.67 11.53$\pm$ 0.91 400.50$\pm$ 5.36 4.45$\pm$ 0.39
UwV83HsGsw3 28.10$\pm$ 0.67 15.16$\pm$ 2.66 398.37$\pm$ 9.11 4.95$\pm$ 0.19
Vt2qJdWjCF2 18.33$\pm$ 3.12 7.74$\pm$ 1.46 443.81$\pm$ 10.96 11.66$\pm$ 0.67
VzqfbhrpDEA 30.00$\pm$ 0.91 16.90$\pm$ 0.92 387.73$\pm$ 5.64 7.28$\pm$ 0.43
X7HyMhZNoso 63.33$\pm$ 0.89 32.24$\pm$ 0.69 270.01$\pm$ 2.15 2.31$\pm$ 0.39
Z6MFQCViBuw 40.67$\pm$ 3.40 15.79$\pm$ 3.40 380.14$\pm$ 11.92 4.87$\pm$ 0.44
gTV8FGcVJC9 46.50$\pm$ 1.50 28.97$\pm$ 1.96 313.77$\pm$ 0.72 5.54$\pm$ 0.94
gYvKGZ5eRqb 10.00 2.73$\pm$ 0.50 471.45$\pm$ 3.05 9.34$\pm$ 1.00
oLBMNvg9in8 47.00$\pm$ 1.00 24.69$\pm$ 1.35 332.24$\pm$ 4.97 1.92$\pm$ 0.16
pLe4wQe7qrG 65.00$\pm$ 1.67 28.04$\pm$ 5.24 290.87$\pm$ 27.90 0.58$\pm$ 0.09
pa4otMbVnkk 33.33$\pm$ 3.33 18.79$\pm$ 0.48 365.55$\pm$ 13.95 7.17$\pm$ 0.66
rqfALeAoiTq 43.89$\pm$ 2.19 22.01$\pm$ 1.75 348.34$\pm$ 8.51 2.55$\pm$ 0.29
x8F5xyUWy9e 52.42$\pm$ 0.43 22.44$\pm$ 0.94 328.36$\pm$ 2.28 2.13$\pm$ 0.02
yqstnuAEVhm 38.61$\pm$ 3.14 17.06$\pm$ 2.82 372.31$\pm$ 14.27 2.83$\pm$ 0.13
zsNo4HB9uLZ 26.67$\pm$ 1.18 16.10$\pm$ 1.78 400.21$\pm$ 5.77 4.74$\pm$ 0.08
Average 37.93 18.80 368.01 4.72
Table 24: NOLO(M) performance in Habitat testing scenes.
[…]
2azQ1b91cZZ 35.77$\pm$ 0.38 21.70$\pm$ 0.98 367.77$\pm$ 0.44 5.26$\pm$ 0.06
8194nk5LbLH 38.18$\pm$ 0.91 19.20$\pm$ 0.74 365.92$\pm$ 3.05 3.05$\pm$ 0.18
EU6Fwq7SyZv 40.50$\pm$ 4.50 20.97$\pm$ 1.42 360.27$\pm$ 14.82 3.04$\pm$ 0.49
QUCTc6BB5sX 40.00 22.11$\pm$ 1.36 354.65$\pm$ 1.51 4.64$\pm$ 0.25
SN83YJsR3w2 32.50$\pm$ 7.50 10.28$\pm$ 2.55 414.92$\pm$ 23.22 4.92$\pm$ 1.20
TbHJrupSAjP 24.29$\pm$ 1.43 7.48$\pm$ 0.50 427.30$\pm$ 6.51 4.63$\pm$ 0.07
UwV83HsGsw3 30.71$\pm$ 2.14 17.89$\pm$ 3.05 390.06$\pm$ 15.81 4.03$\pm$ 0.21
Vt2qJdWjCF2 13.75$\pm$ 1.25 6.65$\pm$ 0.62 454.84$\pm$ 0.97 12.59$\pm$ 0.18
VzqfbhrpDEA 28.33$\pm$ 3.89 14.89$\pm$ 2.38 402.32$\pm$ 9.62 7.08$\pm$ 0.47
X7HyMhZNoso 64.64$\pm$ 1.79 30.95$\pm$ 0.11 276.48$\pm$ 5.50 1.94$\pm$ 0.15
Z6MFQCViBuw 37.00$\pm$ 7.00 16.86$\pm$ 3.37 374.19$\pm$ 22.49 4.90$\pm$ 0.33
gTV8FGcVJC9 50.50$\pm$ 7.50 28.67$\pm$ 1.23 306.25$\pm$ 9.88 4.54$\pm$ 0.72
gYvKGZ5eRqb 12.50$\pm$ 2.50 3.79$\pm$ 1.73 475.62$\pm$ 14.77 9.55$\pm$ 1.20
oLBMNvg9in8 48.00$\pm$ 2.00 28.67$\pm$ 0.73 319.90$\pm$ 6.13 1.61$\pm$ 0.16
pLe4wQe7qrG 53.33$\pm$ 3.33 26.82$\pm$ 1.01 311.75$\pm$ 0.38 1.26$\pm$ 0.19
pa4otMbVnkk 20.00 8.13$\pm$ 0.40 431.58$\pm$ 1.42 8.87$\pm$ 0.33
rqfALeAoiTq 45.42$\pm$ 2.08 20.42$\pm$ 2.06 352.25$\pm$ 7.54 2.19$\pm$ 0.23
x8F5xyUWy9e 49.55$\pm$ 4.09 22.50$\pm$ 1.80 331.82$\pm$ 12.30 2.14$\pm$ 0.47
yqstnuAEVhm 42.08$\pm$ 4.58 18.98$\pm$ 3.12 360.02$\pm$ 16.95 2.89$\pm$ 0.05
zsNo4HB9uLZ 27.50 16.36$\pm$ 1.47 400.51$\pm$ 1.49 4.76$\pm$ 0.10
Average 36.73 18.17 373.92 4.69
9.3 Cross Evaluation Results
The performances of H2R and R2H are reported in Tabs. 25 and 26 respectively.
Table 25: NOLO performance in RoboTHOR after training in Habitat. “T” denotes FloorPlan_Train and “V” denotes FloorPlan_Val.
[…]
T10_5 60.00$\pm$ 0.96 26.63$\pm$ 1.37 303.93$\pm$ 7.01 2.30$\pm$ 0.13
T11_5 64.89$\pm$ 3.50 31.85$\pm$ 3.64 286.70$\pm$ 20.00 2.09$\pm$ 0.22
T12_5 52.16$\pm$ 5.79 26.13$\pm$ 3.37 326.15$\pm$ 15.30 2.36$\pm$ 0.24
T1_5 50.00$\pm$ 6.24 29.40$\pm$ 1.07 320.91$\pm$ 10.47 2.46$\pm$ 0.04
T2_5 47.94$\pm$ 4.56 20.57$\pm$ 4.91 339.62$\pm$ 26.92 2.42$\pm$ 0.18
T3_5 39.22$\pm$ 3.09 18.48$\pm$ 0.60 377.95$\pm$ 6.26 3.30$\pm$ 0.22
T4_5 64.24$\pm$ 1.87 30.56$\pm$ 0.79 279.33$\pm$ 3.76 2.13$\pm$ 0.13
T5_5 61.52$\pm$ 0.86 27.48$\pm$ 2.16 305.47$\pm$ 7.33 2.06$\pm$ 0.11
T6_5 60.56$\pm$ 2.83 31.06$\pm$ 4.18 285.57$\pm$ 25.83 2.30$\pm$ 0.16
T7_5 50.83$\pm$ 0.59 20.54$\pm$ 1.48 334.78$\pm$ 0.93 2.30$\pm$ 0.17
T8_5 54.58$\pm$ 4.12 28.68$\pm$ 1.69 297.57$\pm$ 10.62 2.41$\pm$ 0.02
T9_5 38.67$\pm$ 6.60 22.46$\pm$ 4.39 364.06$\pm$ 18.55 2.53$\pm$ 0.15
Average 53.72 26.15 318.50 2.39
V1_1 54.00$\pm$ 4.32 28.45$\pm$ 3.73 316.52$\pm$ 19.28 2.34$\pm$ 0.15
V1_2 48.48$\pm$ 4.09 23.46$\pm$ 3.96 338.57$\pm$ 19.55 2.61$\pm$ 0.12
V1_3 74.04$\pm$ 4.06 39.60$\pm$ 2.33 235.18$\pm$ 8.91 1.71$\pm$ 0.09
V1_4 44.07$\pm$ 5.83 23.73$\pm$ 1.69 350.40$\pm$ 17.36 2.65$\pm$ 0.12
V1_5 49.44$\pm$ 4.10 27.90$\pm$ 4.66 324.46$\pm$ 26.83 2.68$\pm$ 0.17
V2_1 32.73$\pm$ 3.71 14.01$\pm$ 1.79 404.18$\pm$ 12.65 3.19$\pm$ 0.10
V2_2 48.89$\pm$ 0.91 24.29$\pm$ 3.87 338.47$\pm$ 12.65 2.96$\pm$ 0.19
V2_3 67.18$\pm$ 1.92 35.17$\pm$ 2.13 278.73$\pm$ 10.96 1.63$\pm$ 0.08
V2_4 52.69$\pm$ 0.38 25.33$\pm$ 0.32 340.45$\pm$ 7.34 2.33$\pm$ 0.03
V2_5 44.35$\pm$ 2.61 22.94$\pm$ 1.82 353.16$\pm$ 22.72 2.94$\pm$ 0.11
V3_1 59.05$\pm$ 2.86 29.58$\pm$ 2.96 279.42$\pm$ 18.76 2.24$\pm$ 0.23
V3_2 56.39$\pm$ 1.71 22.96$\pm$ 1.40 310.36$\pm$ 17.14 2.40$\pm$ 0.08
V3_3 59.51$\pm$ 4.58 26.33$\pm$ 3.53 297.91$\pm$ 19.58 2.22$\pm$ 0.14
V3_4 41.59$\pm$ 1.19 15.47$\pm$ 1.16 360.06$\pm$ 9.11 2.95$\pm$ 0.01
V3_5 60.00$\pm$ 2.36 27.78$\pm$ 2.98 298.36$\pm$ 11.92 2.15$\pm$ 0.15
Average 52.83 25.80 321.75 2.47
Table 26: NOLO Performance in Habitat testing scenes after training in RoboTHOR.
[…]
2azQ1b91cZZ 38.21$\pm$ 0.96 23.38$\pm$ 0.54 355.25$\pm$ 4.81 6.45$\pm$ 0.16
8194nk5LbLH 41.82$\pm$ 2.23 21.25$\pm$ 0.96 351.39$\pm$ 6.97 4.85$\pm$ 0.14
EU6Fwq7SyZv 45.67$\pm$ 2.87 22.27$\pm$ 1.75 345.99$\pm$ 10.11 4.72$\pm$ 0.37
QUCTc6BB5sX 44.00$\pm$ 2.45 23.97$\pm$ 1.79 346.11$\pm$ 11.27 6.91$\pm$ 0.19
SN83YJsR3w2 18.33$\pm$ 4.71 7.28$\pm$ 0.33 449.10$\pm$ 9.22 9.20$\pm$ 0.77
TbHJrupSAjP 28.10$\pm$ 2.94 8.99$\pm$ 0.80 422.23$\pm$ 3.90 6.39$\pm$ 0.26
UwV83HsGsw3 27.62$\pm$ 4.10 14.83$\pm$ 3.12 408.73$\pm$ 7.70 7.07$\pm$ 0.60
Vt2qJdWjCF2 17.92$\pm$ 2.12 8.39$\pm$ 1.65 436.55$\pm$ 5.17 14.59$\pm$ 0.41
VzqfbhrpDEA 27.41$\pm$ 2.28 15.64$\pm$ 0.72 396.57$\pm$ 6.72 9.19$\pm$ 0.40
X7HyMhZNoso 62.86$\pm$ 1.17 29.77$\pm$ 1.18 278.76$\pm$ 4.89 4.16$\pm$ 0.42
Z6MFQCViBuw 47.33$\pm$ 3.40 15.48$\pm$ 1.83 368.63$\pm$ 8.25 5.93$\pm$ 0.66
gTV8FGcVJC9 50.00$\pm$ 1.63 27.88$\pm$ 2.24 316.60$\pm$ 11.00 7.03$\pm$ 0.44
gYvKGZ5eRqb 18.33$\pm$ 2.36 3.68$\pm$ 0.50 468.87$\pm$ 6.26 9.14$\pm$ 1.13
oLBMNvg9in8 47.67$\pm$ 1.70 27.95$\pm$ 2.42 322.73$\pm$ 10.51 3.65$\pm$ 0.08
pLe4wQe7qrG 51.11$\pm$ 1.57 26.30$\pm$ 4.24 317.16$\pm$ 15.91 3.56$\pm$ 0.22
pa4otMbVnkk 31.11$\pm$ 9.56 13.82$\pm$ 3.98 398.71$\pm$ 26.33 9.85$\pm$ 1.45
rqfALeAoiTq 38.61$\pm$ 3.22 18.14$\pm$ 1.66 369.56$\pm$ 11.93 4.89$\pm$ 0.11
x8F5xyUWy9e 50.00$\pm$ 3.40 21.01$\pm$ 1.50 332.68$\pm$ 12.40 4.41$\pm$ 0.13
yqstnuAEVhm 41.11$\pm$ 2.39 18.79$\pm$ 1.28 363.67$\pm$ 11.33 4.81$\pm$ 0.08
zsNo4HB9uLZ 28.75$\pm$ 2.70 16.22$\pm$ 0.45 399.03$\pm$ 8.76 6.50$\pm$ 0.24
Average 37.80 18.25 372.41 6.67
10 Social Impact and Future Work
We see our work as a starting point to propose a general and scalable method to offline learn a navigation policy from limited sources for any mobile agents. In the future, We believe NOLO will benefit a lot from the rapid development in optical flow and computer vision foundation models, and there will be extensive applications of NOLO, such as deployment to sweeping robots or rescue robots. Further improvement can be conducted to examine and enhance the reliability of the decoded actions from optical flow. It is also a great idea to incorporate first-person and third-person videos to showcase more advanced intelligence of spatial understanding and egocentric decision-making in more complex 3D embodied environments.